OpenStudy (anonymous):

The length of a rectangular field is 4 m longer than its width. If the perimeter is 56 m, what is the width of the field? A. 8 m B. 12 m C. 14 m D. 16 m

OpenStudy (jordanloveangel):

@Clonetroooper78 u still need help cuz i got ur answer?

OpenStudy (anonymous):

l=4+w

OpenStudy (anonymous):

and 2l+2w=56

OpenStudy (jordanloveangel):

Let x = width then, length= x+4 2 * (width + length) = perimeter 2*(x +(x + 4)) = 56 2*( 2x+4) = 56 4x + 8 = 56 4x = 48 x = 12m

OpenStudy (jordanloveangel):

so ur answer is 12m

OpenStudy (anonymous):

so you get 12 B

OpenStudy (jordanloveangel):

yup